Give the coordinates of two points on a coordinate grid that are square root of 10 apart.
Answer by richwmiller(17219)

Working backwards with the distance formula to find the points
sqrt(10)^2=10
a^2+b^2=10
1^2+3^2=10
(1-0)^2+(3-0)^2)
(0,0) and (3,1)
There are of course many more combinations.